OpenStudy (anonymous):

Solve. u + 6 = 18 A. 3 B. 12 C. 24 D. 108

OpenStudy (anonymous):

please help

OpenStudy (youngstudier):

Ok so we need to solve for u, right? Ok, so, first we need to subtract 6 from both sides of the equation.

OpenStudy (anonymous):

ok

OpenStudy (youngstudier):

So it would look like this: u + 6 - 6 = 18 -6

OpenStudy (youngstudier):

It'll get rid of 6 on the u side. So, what's 18 -6?

OpenStudy (anonymous):

12

OpenStudy (youngstudier):

Yup, so: u = 12. Good job!

OpenStudy (anonymous):

thanks

OpenStudy (youngstudier):

No problem! ^.^

OpenStudy (ahvei):

B is the answer

OpenStudy (anonymous):

can you help me with somemore

OpenStudy (ahvei):

who me?

OpenStudy (youngstudier):

Sure! :D

OpenStudy (anonymous):

Solve. m – 14.6 = 9.74 A. 4.86 B. 23.34 C. 24.34 D. 25.34

OpenStudy (youngstudier):

ok so first, we need to add 14.6 to both sides of the equation, right?

OpenStudy (anonymous):

yeah

OpenStudy (youngstudier):

So it would look like this: m - 14.6 + 14.6 = 9.74 + 14.6

OpenStudy (ahvei):

the answer is C because if you add them up you will get the answer

OpenStudy (ahvei):

always check ur answer tho sumtimes it best to plug in all the answers

OpenStudy (anonymous):

thanks

OpenStudy (ahvei):

your welcome have anymore

OpenStudy (anonymous):

yes

OpenStudy (youngstudier):

@ahvei , please don't give her the answer. I'm trying to help her get the answer.

OpenStudy (youngstudier):

By herself.

OpenStudy (ahvei):

i did both

OpenStudy (anonymous):

Which equation can be used to find the answer? A playground with four sides has a perimeter of 73 ft. Three of the sides have lengths of 14 ft, 19 ft, and 25 ft. What is the length of the fourth side? A. s + 14 + 19 + 25 = 73 s = 15 15 ft B. s – 14 + 19 + 25 = 73 s =43 43 ft C. s – 73 = 14 + 19 – 25 s = 81 81 ft D. s – 14 – 19 – 25 = 73 s = 131 131 ft

OpenStudy (youngstudier):

This way she can better understand it so she won't get stuck again.

OpenStudy (ahvei):

A is the answer ok if you add all of them up you will get 58 so when u subtract 73 by 58 you get 15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i have one more

OpenStudy (ahvei):

anymore...if you want me to explain it better i can

OpenStudy (anonymous):

Find the value of the unknown side length for the given perimeter P. P = 14 m A. 3 m B. 4 m C. 14 m D. 25 m

OpenStudy (anonymous):

here is the last question

OpenStudy (ahvei):

D is the answer...because if you add them all up and since uk what the varible is worth you add that up too and you would get 25 as the perimeter

OpenStudy (anonymous):

thank you so much

OpenStudy (ahvei):

you are very welcome come to me anyday if you need help

OpenStudy (anonymous):